# Energy, enstrophy and helicity transfers in polymeric turbulence

### v1

@@ 2.2 The budget equations
The various terms appearing in the expressions of the fluxes are interpreted as the mean and the turbulent transport, pressure transport and viscous diffusion. Note that pressure does not play a role for $\bm{\psi}^{\delta\omega^{2}}$ due to the curl operator. As already stated in Baj et al. (2022), the budget equation for $\left\langle{\delta h}\right\rangle$ shows that the transfer of helicity in the $\bm{X}$ and $\bm{r}$ spaces associated with turbulence fluctuations arise not only from the interaction of $\delta h$ with the velocity increments $\delta u_{j}$, but also from the interaction of $\delta q^{2}$ with the vorticity increments $\delta\omega_{j}$. Accordingly, there result two different routes of helicity transfer, similar to the results of Yan et al. (2020). The presence of polymers introduces an additional contribution to $\bm{\psi}$ and $\bm{\phi}$, identified as polymeric transport. The right hand side of equation 4 describes the net source of $\left\langle{\delta q^{2}}\right\rangle$, $\left\langle{\delta\omega^{2}}\right\rangle$ and $\left\langle{\delta h}\right\rangle$ in space and among scales. The mean production terms identify how $\left\langle{\delta q^{2}}\right\rangle$, $\left\langle{\delta\omega^{2}}\right\rangle$ and $\left\langle{\delta h}\right\rangle$ are exchanged with the mean flow. Accordingly, similar to the single-point budget equations, the production terms are related with the mean-flow gradients. Unlike for $\left\langle{\delta q^{2}}\right\rangle$, the mean production of $\left\langle{\delta h}\right\rangle$ and $\left\langle{\delta\omega^{2}}\right\rangle$ depends on both $\partial U_{i}/\partial x_{j}$ and $\partial\Omega_{i}/\partial x_{j}$. Likewise, enstrophy production receives additional contributions due to vortex stretching by turbulent fluctuations. When present, the mean vorticity gradient also contributes to vortex stretching, and therefore to the production of $\left\langle{\delta h}\right\rangle$ and $\left\langle{\delta\omega^{2}}\right\rangle$. In addition to the fluid dissipation terms $\mathcal{E}$, the right hand side of equation 4 features an additional sink/source $\Pi$ which describes the exchange of $\left\langle{\delta q^{2}}\right\rangle$, $\left\langle{\delta\omega^{2}}\right\rangle$ and $\left\langle{\delta h}\right\rangle$ with the polymeric microstructure. This term is not positive-definite, meaning that exchanges in either direction are possible locally.
@@ 2.2.1 Homogeneous isotropic turbulence
For homogeneous flows, the three structure functions and the terms of the corresponding budget equations lose their dependence on $\bm{X}$, and are functions of only $\bm{r}\equiv(r_{x},r_{y},r_{z})$. Exploiting isotropy, we integrate in the $\bm{r}$ space over spherical shells of radius $r=|\bm{r}|$, with surface $S(r)$ and volume $\Omega(r)$, and obtain the generalised Kármán-Howarth equation for homogeneous isotropic polymeric turbulence. The three relations respectively for the velocity ($\left\langle{\delta q^{2}}\right\rangle$), vorticity ($\left\langle{\delta\omega^{2}}\right\rangle$) and helicity ($\left\langle{\delta h}\right\rangle$) structure functions are given as follows, i.e.
$\begin{gathered}\underbrace{\frac{1}{S(r)}\oint_{S(r)}\left\langle{\delta u_{j}\delta q^{2}}\right\rangle n_{j}\text{d}S}_{\left\langle{\Phi_{f}^{\delta q^{2}}}\right\rangle(r)}+\underbrace{\frac{1}{S(r)}\oint_{S(r)}-\frac{4}{\rho}\left\langle{\delta u_{i}T_{ij}^{*}}\right\rangle n_{j}\text{d}S}_{\left\langle{\Phi_{p}^{\delta q^{2}}}\right\rangle(r)}=\underbrace{\frac{1}{S(r)}\oint_{S(r)}2\nu\frac{\partial\left\langle{\delta q^{2}}\right\rangle}{\partial r_{j}}n_{j}\text{d}S}_{\left\langle{D_{f}^{\delta q^{2}}}\right\rangle(r)}\\ -\frac{4}{3}\underbrace{\nu\left\langle{\frac{\partial u_{i}}{\partial x_{j}}\frac{\partial u_{i}}{\partial x_{j}}}\right\rangle}_{\left\langle{\varepsilon_{f}^{\delta q^{2}}}\right\rangle}r-\frac{4}{3}\underbrace{\frac{1}{\rho}\left\langle{T_{ij}\frac{\partial u_{i}}{\partial x_{j}}}\right\rangle}_{\left\langle{\pi^{\delta q^{2}}}\right\rangle}r+\underbrace{\frac{1}{S(r)}\int_{V(r)}2\left\langle{\delta u_{i}\delta f_{i}}\right\rangle\text{d}V}_{\left\langle{F^{\delta q^{2}}}\right\rangle(r)},\end{gathered}$ (17)
$\begin{gathered}\underbrace{\frac{1}{S(r)}\oint_{S(r)}\left\langle{\delta u_{j}\delta\omega^{2}}\right\rangle n_{j}\text{d}S}_{\left\langle{\Phi_{f}^{\delta\omega^{2}}}\right\rangle(r)}+\underbrace{\frac{1}{S(r)}\oint_{S(r)}-\frac{4}{\rho}\left\langle{\delta\omega_{i}T^{\omega*}_{ij}}\right\rangle n_{j}\text{d}S}_{\left\langle{\Phi_{p}^{\delta\omega^{2}}}\right\rangle(r)}=\underbrace{\frac{1}{S(r)}\oint_{S(r)}2\nu\frac{\partial\left\langle{\delta\omega^{2}}\right\rangle}{\partial r_{j}}n_{j}\text{d}S}_{\left\langle{D_{f}^{\delta\omega^{2}}}\right\rangle(r)}\\ +\underbrace{\frac{1}{S(r)}\int_{V(r)}2\left(\left\langle{\delta\omega_{i}\omega_{j}^{*}\delta\left(\frac{\partial u_{i}}{\partial x_{j}}\right)}\right\rangle+\left\langle{\delta\omega_{i}\delta\omega_{j}\left(\frac{\partial u_{i}}{\partial x_{j}}\right)^{*}}\right\rangle\right)\text{d}V}_{\left\langle{V_{st}^{\delta\omega^{2}}}\right\rangle(r)}-\frac{4}{3}\underbrace{\nu\left\langle{\frac{\partial\omega_{i}}{\partial x_{j}}\frac{\partial\omega_{i}}{\partial x_{j}}}\right\rangle}_{\left\langle{\varepsilon_{f}^{\delta\omega^{2}}}\right\rangle}r\\ -\frac{4}{3}\underbrace{\frac{1}{\rho}\left\langle{T^{\omega}_{ij}\frac{\partial\omega_{i}}{\partial x_{j}}}\right\rangle}_{\left\langle{\pi^{\delta\omega^{2}}}\right\rangle}r+\underbrace{\frac{1}{S}(r)\int_{V(r)}2\left\langle{\delta f^{\omega}_{i}\delta\omega_{i}}\right\rangle\text{d}V}_{\left\langle{F^{\delta\omega^{2}}}\right\rangle(r)}\end{gathered}$ (18)
@@
$\begin{gathered}\underbrace{\frac{1}{S(r)}\oint_{S(r)}\left\langle{\delta u_{j}\delta h}\right\rangle n_{j}\text{d}S}_{\left\langle{\Phi_{f,a}^{\delta h}}\right\rangle(r)}+\underbrace{\frac{1}{S(r)}\oint_{S(r)}-\frac{1}{2}\left\langle{\delta\omega_{j}\delta q^{2}}\right\rangle n_{j}\text{d}S}_{\left\langle{\Phi_{f,b}^{\delta h}}\right\rangle(r)}\\ +\underbrace{\frac{1}{S(r)}\oint_{S(r)}-\frac{2}{\rho}\left(\left\langle{T_{ij}^{\omega*}\delta u_{i}}\right\rangle+\left\langle{T_{ij}^{*}\delta\omega_{i}}\right\rangle\right)n_{j}\text{d}S}_{\left\langle{\Phi_{p}^{\delta h}}\right\rangle(r)}=\underbrace{\frac{1}{S(r)}\oint_{S(r)}2\nu\frac{\partial\left\langle{\delta h}\right\rangle}{\partial r_{j}}n_{j}\text{d}S}_{\left\langle{D_{f}^{\delta h}}\right\rangle(r)}\\ -\frac{4}{3}\underbrace{\nu\left\langle{\frac{\partial u_{i}}{\partial x_{j}}\frac{\partial\omega_{i}}{\partial x_{j}}}\right\rangle}_{\left\langle{\varepsilon_{f}^{\delta h}}\right\rangle}r-\frac{4}{3}\underbrace{\left\langle{\frac{1}{2\rho}\left(T_{ij}^{\omega}\frac{\partial u_{i}}{\partial x_{j}}+T_{ij}\frac{\partial\omega_{i}}{\partial x_{j}}\right)}\right\rangle}_{\left\langle{\pi^{\delta h}}\right\rangle}r+\underbrace{\frac{1}{S(r)}\int_{V(r)}\left(\left\langle{\delta f^{\omega}_{i}\delta u_{i}}\right\rangle+\left\langle{\delta f_{i}\delta\omega_{i}}\right\rangle\right)\text{d}V}_{\left\langle{F^{\delta h}}\right\rangle(r)}.\end{gathered}$ (19)
@@
These equations reveal the main differences between Newtonian and polymeric turbulence: for Newtonian fluids, energy is dissipated by only the fluid viscosity $\varepsilon_{f}^{\delta q^{2}}$, and the flux of energy is only due to the nonlinear term $\Phi_{f}^{\delta q^{2}}$, alongside the assumption that in the inertial range of scales $D_{f}^{\delta q^{2}}$ and $F^{\delta q^{2}}$ are negligible. In this range, a (positive) production term associated with vortex stretching arises only for the enstrophy equation. For polymeric flows, the total flux comprises an additional component $\Phi_{p}$ that mimics the energy transfer among scales associated with the polymeric microstructure. An additional source/sink term $\pi$ appears on the right hand side, that characterises the exchange of energy, helicity and enstrophy between the fluid and the polymers. It is interesting to note that for the energy, $\langle\pi^{\delta q^{2}}\rangle$ matches the polymeric dissipation $\langle\epsilon_{p}\rangle\equiv\left\langle{\mu_{p}(R_{ii}-3)/(2\tau_{p}^{2})}\right\rangle$, meaning that on average this term is a net sink for the fluid phase (De Angelis et al., 2005).
@@
When studying the scale-by-scale energy budget of polymeric turbulence in Fourier space, Rosti et al. (2023) showed that the non Newtonian flux is not a purely dissipative term. They isolate the purely dissipative part of this flux by assuming that it has the same asymptotic dependence as the fluid viscous flux at small scales. An additional requirement is that at vanishing scales it matches the polymeric dissipation $\langle\pi^{\delta q^{2}}\rangle=\langle\varepsilon_{p}\rangle$. Here we use the same approach for the three budgets. We thus split the polymeric flux $\Phi_{p}$ into two contributions,
@@ 2.3 The numerical database
This work relies on the DNS database introduced by Singh & Rosti (2024). Their paper contains full details on the numerical method and the related computational procedures, which are only briefly recalled here. Equations 1 are numerically integrated using the in-house solver Fujin (https://www.oist.jp/research/research-units/cffu/fujin), which uses an incremental pressure-correction scheme. The Navier–Stokes equations written in primitive variables are solved on a staggered grid using second-order finite-differences in all the directions. The momentum equation is advanced in time using a second-order Adams-Bashforth time scheme, while the non Newtonian stress tensor is advanced in time with a second-order Crank-Nicolson scheme. A log-conformation formulation (Izbassarov et al., 2018) ensures positive-definiteness of the conformation tensor at all times. Turbulence is sustained using the Arnold-Beltrami-Childress (ABC) cellular-flow forcing (Podvigina & Pouquet, 1994).
@@
The equations are solved within a cubic domain of size $L=2\pi$ having periodic boundary conditions in all directions, discretised with $N^{3}=1024^{3}$ grid points, to ensure that all the scales down to the smallest dissipative ones are properly resolved, i.e. $\eta/\Delta x=O(1)$, where $\Delta x$ is the grid spacing and $\eta$ the Kolmogorov scale. The parameters are chosen to achieve a Taylor-microscale Reynolds number of $Re_{\lambda}=u^{\prime}\lambda/\nu\approx 460$ in the purely Newtonian case ($u^{\prime}$ is the root mean square of the velocity fluctuations and $\lambda$ is the Taylor length scale). The Deborah number $De=\tau_{p}/\tau_{f}$, where $\tau_{f}=L/u_{rms}$ is the large-eddy turnover time, is varied in the range $1/9\leq De\leq 9$. For all cases, the fluid and polymer viscosities are fixed such that $\mu_{f}/(\mu_{f}+\mu_{p})=0.9$. Simulations are advanced with a constant time step of $\Delta t/\tau_{\eta}=2\times 10^{-3}$, where $\tau_{\eta}$ is the Kolmogorov time scale. Details of the numerical simulations with bulk quantities of interest are provided in table 1.
@@
Table 1: Details of the numerical simulations considered in the present study. $De$ is the Deborah number, $Re_{\lambda}$ is the Reynolds number based on $u^{\prime}=\sqrt{2E/3}$ and on the Taylor length scale $\lambda$, $\eta$ is the Kolmogorov length scale, $R_{ii}$ is the trace of the conformation tensor and mimics the free energy of the polymeric phase, $\varepsilon_{f}^{\delta q^{2}}$, $\varepsilon_{f}^{\delta\omega^{2}}$ and $\varepsilon_{f}^{\delta h}$ are the fluid dissipation of energy, enstrophy and helicity, and $\pi^{\delta q^{2}}$, $\pi^{\delta\omega^{2}}$, $\pi^{\delta h}$ are the source/sink of energy, enstrophy and helicity due to the fluid-polymer coupling.
@@
$De$ $Re_{\lambda}$ $\left\langle{\lambda}\right\rangle$ $\left\langle{\eta}\right\rangle$ $\langle\varepsilon_{f}^{\delta q^{2}}\rangle$ $\langle\varepsilon_{f}^{\delta\omega^{2}}\rangle$ $\langle\varepsilon_{f}^{\delta h^{2}}\rangle$ $\langle\pi^{\delta q^{2}}\rangle$ $\langle\pi^{\delta\omega^{2}}\rangle$ $\langle\pi^{\delta h}\rangle$
-- $458.6$ $0.1578$ $0.0037$ $59.82$ $5.1041\times 10^{5}$ $61.51$- -- --
$1/9$ $513.64$ $0.1747$ $0.0039$ $50.84$ $3.6950\times 10^{5}$ $55.14$ $6.50$ $2.5279\times 10^{4}$ $5.51$
$1/3$ $627.06$ $0.2037$ $0.0041$ $41.76$ $2.4098\times 10^{5}$ $44.38$ $17.29$ $2.0524\times 10^{4}$ $12.76$
$1$ $795.80$ $0.2767$ $0.0050$ $21.04$ $1.1265\times 10^{5}$ $33.09$ $39.94$ $-7.0836\times 10^{4}$ $33.84$
$3$ $761.11$ $0.2859$ $0.0053$ $15.50$ $9.1522\times 10^{4}$ $25.21$ $33.52$ $-7.7998\times 10^{4}$ $22.48$
$9$ $776.30$ $0.2733$ $0.0050$ $22.03$ $1.2596\times 10^{5}$ $33.06$ $45.61$ $-1.0975\times 10^{5}$ $36.42$

### v2

@@ 2.2 The budget equations
The various terms in the fluxes $\{\bm{\phi},\bm{\psi}\}$ can be interpreted as contributions from mean and turbulent transport and pressure transport. Notably, the pressure does not contribute to $\bm{\psi}^{\delta\omega^{2}}$, as its effect is eliminated by the curl operator. In addition to the standard nonlinear transfer terms, the fluxes include an additional contribution that reflects a distinct transfer mechanism induced by fluid-polymer interactions, which is absent in purely Newtonian flows. As previously noted by Baj et al. (2022), the budget equation for $\langle\delta h\rangle$ reveals that helicity transfer in both physical space ($\bm{X}$) and scale space ($\bm{r}$) arises not only from the interaction between $\delta h$ and the velocity increments $\delta u_{j}$, but also from the coupling between $\delta q^{2}$ and the vorticity increments $\delta\omega_{j}$. These two distinct mechanisms give rise to separate pathways for helicity transfer, consistent with the findings of Yan et al. (2020). The right-hand side of equation (6) represents the net source terms for $\langle\delta q^{2}\rangle$, $\langle\delta\omega^{2}\rangle$, and $\langle\delta h\rangle$, accounting for exchanges in both physical space and across scales. The mean production terms $\mathcal{P}$ describe how these quantities are exchanged with the mean flow, analogous to the single-point budget equations. As such, they are directly linked to the gradients of the mean velocity field. Unlike $\langle\delta q^{2}\rangle$, the mean production of $\langle\delta\omega^{2}\rangle$ and $\langle\delta h\rangle$ depends not only on the mean velocity gradients $\partial U_{i}/\partial x_{j}$ but also on the gradients of the mean vorticity $\partial\Omega_{i}/\partial x_{j}$. Moreover, enstrophy production includes additional contributions from vortex stretching induced by turbulent fluctuations $\mathfrak{p}^{\delta\omega^{2}}$. When present, gradients in the mean vorticity field further enhance this mechanism. In addition to the fluid dissipation terms $\mathcal{E}$, the right-hand side of equation (6) includes a polymeric contribution $\Pi$, which reflects the influence of fluid-polymer interactions. Unlike dissipation, this term is not sign-definite meaning that it can act as either a source or a sink depending on the local configuration of the polymer stress and velocity gradients, see §5.
@@ 2.2.1 Homogeneous isotropic turbulence
For homogeneous statistically steady flows, the three structure functions and the terms of the corresponding budget equations lose their dependence on $\bm{X}$ and $t$, and are functions of only the separation vector $\bm{r}\equiv(r_{x},r_{y},r_{z})$. In HIT the mean-flow is absent and $\bm{U}=\bm{0}$ and $\bm{\Omega}=\bm{0}$, implying $\mathcal{P}^{\delta g}=0$. Exploiting isotropy, we integrate equation (6) in the $\bm{r}$ space over spherical shells of radius $r=|\bm{r}|$, with surface $\mathcal{S}(r)$ and volume $\mathcal{V}(r)$, and obtain the generalised Kármán-Howarth equation for PHIT (see Chiarini et al., 2025, for further details on the energy equation), which in a compact form reads
$\langle\Phi_{f}^{\delta g}\rangle(r)+\langle\Phi_{p}^{\delta g}\rangle(r)-\langle D_{f}^{\delta g}\rangle(r)-\langle V_{st}^{\delta g}\rangle(r)-\langle F^{\delta g}\rangle(r)=-\frac{4}{3}\langle\varepsilon_{f}^{\delta g}\rangle r-\frac{4}{3}\langle\pi^{\delta g}\rangle r.$ (29)
On the left-hand side, $\Phi_{f}^{\delta g}$ and $\Phi_{p}^{\delta g}$ represent the contributions of fluid-fluid and fluid-polymer interactions, respectively, to the scale-space transfer of $\delta g$, and are given by:
$\langle\Phi_{f}^{\delta q^{2}}\rangle=$ $\ \frac{1}{\mathcal{S}(r)}\oint_{\mathcal{S}(r)}\left\langle{\delta u_{j}\delta q^{2}}\right\rangle n_{j}\text{d}\Sigma,$ (30)
$\langle\Phi_{f}^{\delta\omega^{2}}\rangle=$ $\ \frac{1}{\mathcal{S}(r)}\oint_{\mathcal{S}(r)}\left\langle{\delta u_{j}\delta\omega^{2}}\right\rangle n_{j}\text{d}\Sigma,$ (31)
$\langle\Phi_{f}^{\delta h}\rangle=$ $\underbrace{\frac{1}{\mathcal{S}(r)}\oint_{\mathcal{S}(r)}\left\langle{\delta u_{j}\delta h}\right\rangle n_{j}\text{d}\Sigma}_{\left\langle{\Phi_{f,a}^{\delta h}}\right\rangle}+\underbrace{\frac{1}{\mathcal{S}(r)}\oint_{\mathcal{S}(r)}-\left\langle{\delta\omega_{j}\delta q^{2}}\right\rangle n_{j}\text{d}\Sigma}_{\left\langle{\Phi_{f,b}^{\delta h}}\right\rangle}$ (32)
@@
$\langle\Phi_{p}^{\delta q^{2}}\rangle=$ $\ \frac{1}{\mathcal{S}(r)}\oint_{\mathcal{S}(r)}-4\left\langle{\delta u_{i}T_{ij}^{*}}\right\rangle/\rho\ n_{j}\text{d}\Sigma,$ (33)
$\langle\Phi_{p}^{\delta\omega^{2}}\rangle=$ $\ \frac{1}{\mathcal{S}(r)}\oint_{\mathcal{S}(r)}-4\left\langle{\delta\omega_{i}T_{ij}^{\omega,*}}\right\rangle/\rho\ n_{j}\text{d}\Sigma,$ (34)
$\langle\Phi_{p}^{\delta h}\rangle=$ $\ \frac{1}{\mathcal{S}(r)}\oint_{\mathcal{S}(r)}-2(\left\langle{\delta u_{i}T_{ij}^{\omega,*}}\right\rangle+\left\langle{\delta\omega_{i}T_{ij}^{*}}\right\rangle)/\rho\ n_{j}\text{d}\Sigma,$ (35)
where $\bm{n}$ is the outward-pointing unit vector normal to the spherical surface.
Instead, $V_{st}$ denotes the vortex stretching term, which contributes to enstrophy generation across all scales ($V_{st}^{\delta q^{2}}=V_{st}^{\delta h}=0$), and is expressed as:
$\left\langle{V_{st}^{\delta\omega^{2}}}\right\rangle=\frac{1}{\mathcal{S}(r)}\int_{\mathcal{V}(r)}2\left(\left\langle{\delta\omega_{i}\omega_{j}^{*}\delta\left(\frac{\partial u_{i}}{\partial x_{j}}\right)}\right\rangle+\left\langle{\delta\omega_{i}\delta\omega_{j}\left(\frac{\partial u_{i}}{\partial x_{j}}\right)^{*}}\right\rangle\right)\text{d}\Omega.$ (36)
Viscous diffusion at scale $r$ is expressed as
$\langle D_{f}^{\delta q^{2}}\rangle=$ $\ \frac{1}{\mathcal{S}(r)}\oint_{\mathcal{S}(r)}2\nu\frac{\partial\left\langle{\delta q^{2}}\right\rangle}{\partial r_{j}}n_{j}\text{d}\Sigma,$ (37)
$\langle D_{f}^{\delta\omega^{2}}\rangle=$ $\ \frac{1}{\mathcal{S}(r)}\oint_{\mathcal{S}(r)}2\nu\frac{\partial\left\langle{\delta\omega^{2}}\right\rangle}{\partial r_{j}}n_{j}\text{d}\Sigma,$ (38)
$\langle D_{f}^{\delta h}\rangle=$ $\ \frac{1}{\mathcal{S}(r)}\oint_{\mathcal{S}(r)}2\nu\frac{\partial\left\langle{\delta h}\right\rangle}{\partial r_{j}}n_{j}\text{d}\Sigma,$ (39)
while the forcing term that accounts for the injection of $\left\langle{\delta q^{2}}\right\rangle$, $\left\langle{\delta\omega^{2}}\right\rangle$ and $\left\langle{\delta h}\right\rangle$ at scale $r$ reads
$\langle F^{\delta q^{2}}\rangle=$ $\ \frac{1}{\mathcal{S}(r)}\int_{\mathcal{V}(r)}2\left\langle{\delta f_{i}\delta u_{i}}\right\rangle\text{d}\Omega,$ (40)
$\langle F^{\delta\omega^{2}}\rangle=$ $\ \frac{1}{\mathcal{S}(r)}\int_{\mathcal{V}(r)}2\left\langle{\delta f_{i}^{\omega}\delta\omega_{i}}\right\rangle\text{d}\Omega,$ (41)
$\langle F^{\delta h}\rangle=$ $\ \frac{1}{\mathcal{S}(r)}\int_{\mathcal{V}(r)}(\left\langle{\delta f_{i}\delta\omega_{i}}\right\rangle+\left\langle{\delta f_{i}^{\omega}\delta u_{i}}\right\rangle)\text{d}\Omega.$ (42)
@@
These equations highlight the fundamental differences between Newtonian and polymeric turbulence. In Newtonian fluids, dissipation occurs solely through fluid viscosity, represented by $\varepsilon_{f}$, and the fluxes arise exclusively from the nonlinear term $\Phi_{f}$. In the intermediate range of scales where the viscous diffusion $D_{f}$ and forcing $F$ terms are typically negligible, a positive production term associated with vortex stretching appears only in the enstrophy equation. In contrast, polymeric flows feature an additional flux component, $\Phi_{p}$, which (also) accounts for the transfer across scales mediated by the polymers. This term appears in equation (6) under a divergence in scale-space ($\partial/\partial r_{j}$), analogous to $\Phi_{f}$, and its net contribution averages to zero over the entire triperiodic domain. Physically, this reflects how stretched polymers facilitate transfers across scales, providing an alternative cascade pathway that ultimately leads to dissipation. Furthermore, the presence of polymers introduces an additional source/sink term, $\pi$, on the right-hand side of the equations. For the energy equation, $\langle\pi^{\delta q^{2}}\rangle$ matches the polymeric dissipation rate $\langle\epsilon_{p}\rangle\equiv\left\langle{\mu_{p}(R_{ii}-3)/(2\tau_{p}^{2})}\right\rangle$, and on average acts as a net sink for the fluid phase (De Angelis et al., 2005). For the enstrophy and the helicity the average $\langle\pi^{\delta\omega^{2}}\rangle$ and $\langle\pi^{\delta h}\rangle$ can be both positive and negative, depending on the small-scale flow topology.
In the case of homogeneous and isotropic flows, the quantities $\mathcal{E}$ and $\Pi$ reduce to the single-point terms $\varepsilon_{f}$ and $\pi$, which are independent of the scale $r$. The corresponding equations can therefore be interpreted as balance relations for the average total dissipation rates. At each scale $r$, the sum of the scale-dependent terms on the left-hand side must balance the scale-independent average total dissipation rate, which acts as an invariant of the system. This interpretation is fully consistent with the classical Kármán–Howarth equation for the second-order velocity structure function in HIT. However, as discussed above, a key distinction in the case of PHIT is that the invariant quantity is no longer the fluid dissipation alone, but the combined contribution $\langle\varepsilon_{f}\rangle+\langle\pi\rangle$ (Chiarini et al., 2025). In this context, it is worth noting that the equation for $\langle\delta q^{2}\rangle$ closely resembles the classical 4/5-4/5 law for Newtonian homogeneous and isotropic turbulence. Specifically, in the absence of polymers, in the high Reynolds number limit ($Re\to\infty$), and for scales in the inertial range ($r/L\to 0$), the energy equation reduces to $1/S(r)\oint_{S(r)}\langle\delta u_{j}\delta q^{2}\rangle n_{j}\text{d}\Sigma=-(4/3)\langle\varepsilon_{f}^{\delta q^{2}}\rangle r$, which, under assumptions of incompressibility and isotropy (Hill, 1997), simplifies further to the well-known result $S_{3}(r)\equiv\langle\left(\left(\bm{u}(\bm{x}+\bm{r},t)-\bm{u}(\bm{x},t)\right)\cdot\bm{r}/r\right)^{3}\rangle=-(4/5)\langle\varepsilon_{f}^{\delta q^{2}}\rangle r$.
@@
To better characterise the influence of the polymers on the transfers of energy and helicity, we decompose $\Phi_{p}$ into an inertial transfer term $\Phi_{p,i}$ and a dissipative term $D_{p}$, following the approach introduced by Rosti et al. (2023) and informed by earlier work by Perlekar et al. (2006). Perlekar et al. (2006) proposed modelling this behaviour via a scale-dependent effective viscosity. Alternatively, Rosti et al. (2023) separated the dissipative part by assuming that it shares the same asymptotic scale dependence as the fluid viscous term $D_{f}$ and requiring that at vanishing scales it matches the polymeric dissipation $\langle\pi^{\delta q^{2}}\rangle\equiv\langle\varepsilon_{p}\rangle$. Here we adopt this second approach for the budgets of energy and helicity. We split the polymeric flux $\Phi_{p}$ into two contributions, such that
@@
We do not apply the same decomposition to the budget for $\langle\delta\omega^{2}\rangle$, as at large $De$ the fluid–polymer interaction yields, on average, a positive enstrophy source term; that is $\langle\pi^{\delta\omega^{2}}\rangle<0$ (see §4.2). Unlike in the cases of energy and helicity, $\langle\pi^{\delta\omega^{2}}\rangle$ cannot be thus interpreted, on average, as a purely dissipative contribution.
@@ 2.3 The numerical database
We use data from direct numerical simulations (DNS) of homogeneous, isotropic turbulence in a purely elastic Oldroyd-B polymeric fluid. The DNS dataset was first presented by Singh & Rosti (2025), where comprehensive details on the numerical methods and computational procedures can be found. Here, these aspects are only briefly summarised. The governing equations are numerically integrated using the in-house solver Fujin (https://www.oist.jp/research/research-units/cffu/fujin), which uses an incremental pressure-correction scheme. The Navier–Stokes equations written in primitive variables are solved on a staggered grid using second-order finite-differences in all the directions. The momentum equation is advanced in time using a second-order Adams-Bashforth time scheme, while the non Newtonian stress tensor is advanced in time with a second-order Crank-Nicolson scheme. A log-conformation formulation (Izbassarov et al., 2018) ensures positive-definiteness of the conformation tensor at all times. Turbulence is sustained using the Arnold-Beltrami-Childress (ABC) cellular-flow forcing (Podvigina & Pouquet, 1994).
@@
The equations are solved within a cubic domain of size $L=2\pi$ having periodic boundary conditions in all directions. The parameters are chosen to achieve a Taylor-microscale Reynolds number of $Re_{\lambda}=u^{\prime}\lambda/\nu\approx 460$ in the purely Newtonian case ($u^{\prime}$ is the root mean square of the velocity fluctuations and $\lambda$ is the Taylor length scale). The Deborah number $De=\tau_{p}/\tau_{f}$, where $\tau_{f}=L/u_{rms}$ is the reference large-eddy turnover time of the flow in the purely Newtonian case, is varied in the range $1/9\leq De\leq 9$. For all cases, the fluid and polymer viscosities are fixed such that $\mu_{f}/(\mu_{f}+\mu_{p})=0.9$. The computational domain is discretised with $N^{3}=1024^{3}$ grid points, to ensure that all the scales down to the smallest dissipative ones are resolved, i.e. $\left\langle{\eta}\right\rangle/\Delta x=\mathcal{O}(1)$, where $\Delta x$ is the grid spacing and $\eta$ the Kolmogorov length scale. We have $0.6\lessapprox\left\langle{\eta}\right\rangle/\Delta x\lessapprox 0.82$ for all cases. Simulations are advanced with a constant time step of $\Delta t/\left\langle{\tau_{\eta}}\right\rangle=2\times 10^{-3}$, where $\tau_{\eta}$ is the Kolmogorov time scale. Details of the numerical simulations with bulk quantities of interest are provided in table 1.
Concerning the energy budget equation, the model-independence of the results has been demonstrated both in Fourier space by Rosti et al. (2023) and in physical space by Chiarini et al. (2025).
@@
Table 1: Details of the numerical simulations considered in the present study. DeDe is the Deborah number, $Re_{\lambda}$ is the Reynolds number based on $u^{\prime}=\sqrt{2E/3}$ and on the Taylor length scale $\lambda$, $\eta$ is the Kolmogorov length scale, $R_{ii}$ is the trace of the conformation tensor and mimics the free energy of the polymeric phase, $\varepsilon_{f}^{\delta q^{2}}$, $\varepsilon_{f}^{\delta\omega^{2}}$ and $\varepsilon_{f}^{\delta h}$ are the fluid dissipation of energy, enstrophy and helicity, and $\pi^{\delta q^{2}}$, $\pi^{\delta\omega^{2}}$, $\pi^{\delta h}$ are the source/sink of energy, enstrophy and helicity due to the fluid-polymer coupling.
@@
DeDe $Re_{\lambda}$ $\left\langle{\lambda}\right\rangle$ $\left\langle{\eta}\right\rangle$ $\langle\varepsilon_{f}^{\delta q^{2}}\rangle$ $\langle\varepsilon_{f}^{\delta\omega^{2}}\rangle$ $\langle\varepsilon_{f}^{\delta h^{2}}\rangle$ $\langle\pi^{\delta q^{2}}\rangle$ $\langle\pi^{\delta\omega^{2}}\rangle$ $\langle\pi^{\delta h}\rangle$
- $458.6$ $0.157$ $0.0037$ $59.82$ $5.104\times 10^{5}$ $61.51$ - - -
$1/9$ $513.6$ $0.174$ $0.0039$ $50.84$ $3.695\times 10^{5}$ $55.14$ $6.50$ $2.527\times 10^{4}$ 5.51\ \ 5.51
$1/3$ $627.0$ $0.203$ $0.0041$ $41.76$ $2.409\times 10^{5}$ $44.38$ $17.29$ $2.052\times 10^{4}$ $12.76$
11 $795.8$ $0.276$ $0.0050$ $21.04$ $1.126\times 10^{5}$ $33.09$ $39.94$ $7.083\times 10^{4}$ $33.84$
33 $761.1$ $0.285$ $0.0053$ $15.50$ $9.152\times 10^{4}$ $25.21$ $33.52$ $7.799\times 10^{4}$ $22.48$
99 $776.3$ $0.273$ $0.0050$ $22.03$ $1.259\times 10^{5}$ $33.06$ $45.61$ $1.097\times 10^{5}$ $36.42$